OpenStudy (anonymous):

Please help! 15degrees (using 45degrees and 30degrees) cos15degrees= a. cos 45° cos 30° + sin 45° sin 30° b. sin 45° cos 30° + cos 45° sin 30° c. cos 45° cos 30° - sin 45° sin 30° d. sin 45° cos 30° - cos 45° sin 30°

jimthompson5910 (jim_thompson5910):

15 = 45 - 30 Then use the identity cos(A-B) = cos(A)cos(B) + sin(A)sin(B)

jimthompson5910 (jim_thompson5910):

in this case, A = 45 and B = 30

OpenStudy (anonymous):

Thanks :)